OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (anonymous):

y=-x-4

OpenStudy (anonymous):

which one would it be x – y = –4 x – y = 4 x + y = –4 x + y = 4

OpenStudy (anonymous):

Rearrange mad scientists' equation

OpenStudy (anonymous):

xNiker can you do this one What is the solution set for |–4x – 4| – 28 = –4?

OpenStudy (anonymous):

What is the solution set for |–4x – 4| – 28 = –4?

OpenStudy (anonymous):

Do you know what the definition of |x| is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

|x| is -x, x < 0 x, x > 0 So the steps to solving it your problem is to first collect everything outside the absolute signs on one side: |-4x-4| = 24 But we know that this absolute function can be split into two: +(-4x-4) = 24, x>0 -(-4x-4) = 24, x<0 Now you just solve those two equations and you get your answers